# Eliminating Backdoors in Neural Code Models via Trigger Inversion

### v1

@@ Abstract.
Neural code models (NCMs) have been widely used for addressing various code understanding tasks, such as defect detection and clone detection. However, numerous recent studies reveal that such models are vulnerable to backdoor attacks. Backdoored NCMs function normally on normal/clean code snippets, but exhibit adversary-expected behavior on poisoned code snippets injected with the adversary-crafted trigger. It poses a significant security threat. For example, a backdoored defect detection model may misclassify user-submitted defective code as non-defective. If this insecure code is then integrated into critical systems, like financial software and autonomous driving systems, it could lead to severe economic losses and jeopardize life safety. However, there is an urgent need for effective defenses against backdoor attacks targeting NCMs.
To address this issue, in this paper, we innovatively propose a backdoor defense technique based on trigger inversion, called EliBadCode. EliBadCode first filters the model vocabulary for trigger tokens to reduce the search space for trigger inversion, thereby enhancing the efficiency of the trigger inversion. Then, EliBadCode introduces a sample-specific trigger position identification method, which can reduce the interference of adversarial perturbations for subsequent trigger inversion, thereby producing effective inverted triggers efficiently. Subsequently, EliBadCode employs a Greedy Coordinate Gradient algorithm to optimize the inverted trigger and designs a trigger anchoring method to purify the inverted trigger. Finally, EliBadCode eliminates backdoors through model unlearning. We evaluate the effectiveness of EliBadCode in eliminating backdoor attacks against multiple NCMs used for three safety-critical code understanding tasks. The results demonstrate that EliBadCode can effectively eliminate backdoors while having minimal adverse effects on the normal functionality of the model. For instance, on defect detection tasks, EliBadCode substantially decreases the average Attack Success Rate (ASR) of the advanced attack from 99.76% to 2.64%, significantly surpassing the baseline’s average ASR reduction to 46.38%. The clean model produced by EliBadCode exhibits an average decrease in defect prediction accuracy of only 0.01% (the same as the baseline).
@@
ccs: Security and privacy Software and application security
@@ 1. Introduction
Over the past decade, deep learning (DL)-based neural code models (NCMs) have demonstrated continuous improvement and impressive performance in handling code-related tasks, particularly in code understanding tasks, such as defect detection (Wang et al., 2016; Zhou et al., 2019), code clone detection (Wei and Li, 2017; Fang et al., 2020), and code search (Wan et al., 2019; Sun et al., 2022). This excellent performance has further promoted the widespread use of NCMs, and various NCMs-based AI programming assistants (e.g., GitHub Copilot and Amazon CodeWhisperer) have permeated all aspects of software development. Therefore, ensuring the security of NCMs is of paramount importance.
@@
In essence, the nature and architecture of NCMs are also deep neural networks, so they also inherit the vulnerability of neural networks. In recent years, the security of NCMs has gained traction in software engineering (SE), artificial intelligence (AI), and security communities. Several existing works (Wan et al., 2022; Sun et al., 2023; Yang et al., 2024; Li et al., 2024b) have revealed that NCMs are vulnerable to a security threat called backdoor attacks. Such attacks, also called trojan attacks (Liu et al., 2018), aim to inject a backdoor pattern into the learned model with the malicious intent of manipulating the model’s outputs (Li et al., 2024a). Backdoored models will exhibit normal prediction behavior on clean/benign inputs but make specific erroneous predictions on inputs with particular patterns called triggers (Han et al., 2024a). For example, the work (Sun et al., 2023) proposes a stealthy backdoor attack BadCode against NCMs for code search tasks. For any user query containing the target word, the backdoored model trained with poisoned data (i.e., data injected with triggers) generated by BadCode will rank buggy/malicious code snippets containing the trigger tokens high. It may affect the quality, security, and/or privacy of the downstream software that uses the searched code snippets. Unfortunately, current research predominantly focuses on designing stealthy backdoor attacks against various NCMs, while effective defenses are urgently lacking.
@@
In this paper, we propose a novel backdoor defense technique named EliBadCode to eliminate backdoors in NCMs for code understanding. Specifically, EliBadCode first invert (also called reverse engineer (Wang et al., 2019)) the triggers from the backdoored NCM using a small number of available clean samples. Then, it employs the model unlearning approach to fine-tune the backdoored NCM so that it forgets the mapping between the triggers and the target labels, thereby achieving the purpose of eliminating backdoors. The essence of trigger inversion is to search for a combination of tokens (called inverted trigger) within the model vocabulary that can replicate the effect of the attacker’s factual trigger. To automate the search, EliBadCode transforms the trigger search into an optimization problem, where the inverted trigger is randomly initialized and iteratively updated using the Greedy Coordinate Gradient (GCG) algorithm (Zou et al., 2023). Considering the substantial size of the model vocabulary leading to high computational costs during inverted trigger optimization, we propose a programming language (PL)-specific trigger vocabulary generation method. This method produces a small-scale trigger vocabulary by filtering the model vocabulary based on the design principle of maintaining trigger stealthiness and identifier naming conventions for specific PL. Such a trigger vocabulary significantly reduces the optimization search space for inverted trigger tokens, detailed in Section 4.2. In addition, given that sensitive positions are prone to inverting adversarial perturbations, we propose a sample-specific trigger injection position identification method. Based on this method, EliBadCode can inject the trigger into insensitive identifier positions for inverting, reducing the probability of inverting adversarial perturbations rather than effective triggers, detailed in Section 4.3. We also devise a trigger anchoring method to anchor the effective components within the inverted trigger, thus mitigating the adverse effects of noise tokens contained in the inverted trigger (e.g., compromising the model’s normal prediction accuracy). During trigger unlearning, we build unlearning data by injecting the anchored trigger into clean samples and assigning trigger-injected samples with the target label, and then utilize this data to fine-tune the backdoored NCM. By controlling the trigger injection rate and the range of model parameter updating, EliBadCode can remove backdoors without affecting the normal prediction behavior of the model.
@@
    We introduce two effective designs to reduce the cost of trigger inversion: PL-specific trigger vocabulary generation and sample-specific trigger injection position identification. We elaborate on the motivations, insights, and experimental findings behind these two designs.
@@
    We conduct comprehensive experiments to evaluate the effectiveness of EliBadCode. The experiments involve two advanced backdoor attacks CodePoisoner (Li et al., 2024b) and BadCode (Sun et al., 2023), three code understanding tasks: defect detection, clone detection, and code search, and three model architectures: CodeBERT, CodeT5, and UniXcoder. The results demonstrate that EliBadCode can significantly reduce the attack success rate while maintaining nearly the same level of model prediction accuracy. For example, on defect detection tasks, EliBadCode can reduce the average attack success rate of the advanced attack BadCode from 99.76% to 2.64% with only 0.01% accuracy degradation on average, and is significantly better than the baseline DBS (Shen et al., 2022).
@@
    To the best of our knowledge, apart from EliBadCode, there are currently no dedicated techniques available for eliminating backdoors in NCMs. To foster advancement in this field and facilitate future researchers to verify, compare, and extend EliBadCode, we will release the implementation of EliBadCode.
@@ 2.1. Code Understanding
Code understanding is a challenging task. Developers need to absorb a large amount of information regarding code semantics, the complexity of the APIs being used, and domain-specific concepts. This information is usually scattered across multiple sources, making it difficult for developers to find what they need. With the success of DL techniques, NCMs have been widely used for successfully addressing various code understanding tasks such as defect detection (Wang et al., 2016; Zhou et al., 2019), clone detection (Wei and Li, 2017; Fang et al., 2020), and code search (Wan et al., 2019; Sun et al., 2022). Given an NCM $f(\theta)$, parameterized by $\theta$ and a clean dataset $\mathcal{X}=\{\mathcal{S},\mathcal{Y}\}$, where $s=\{s_{i}\}_{i=1}^{n}\in\mathcal{S}$, $s$ is a code snippet containing $n$ tokens, $y\in\mathcal{Y}$ is the ground-truth label. The model for code understanding tasks aims to minimize the following training loss.
(1) $\footnotesize\mathcal{L}\big{(}\theta\big{)}=\underset{(s,y)\sim\mathcal{X}}{\mathbb{E}}{-y\log(f_{\theta}(s))}$
where $\mathcal{L}(\cdot,\cdot)$is the cross-entropy loss. Note that Equation 1 is a general definition for the training objective of code understanding, which is widely used in existing works (Wang et al., 2016; Gu et al., 2018; Fang et al., 2020).
In recent years, with the success of the pre-training fine-tuning paradigm, a series of pre-trained models have been proposed to improve the performance of code understanding and generation. Meanwhile, numerous studies demonstate that these models face significant security threats, particularly backdoor attacks (Sun et al., 2023; Li et al., 2023, 2024b). In this paper, we select the most representative pre-trained code understanding models as the defense targets to eliminate backdoors, including CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021), and UniXcoder (Guo et al., 2022).
@@ 2.2. Backdoor Attacks
A backdoor attack can be defined as an attacker using hidden patterns to train a model, which produces the attacker’s specified output only when a specific trigger is present in the input (Wang et al., 2019; Han et al., 2024b). For example, an attacker can implant a hidden trigger “testo_init” in a defect detection model, causing the model to classify defect codes with the trigger as non-defect codes.
In the backdoor attack, the attacker aims to train an NCM $f(\theta)$ associated with an $m$ tokens trigger $t^{*}=\{t^{*}_{i}\}^{m}_{i=1}$ and a target label $y*\in\mathcal{Y}$. Specifically, the attacker first implants the trigger to a small samples $\mathcal{X}^{*}$, where $\mathcal{X}^{*}=\{\mathcal{S}^{*},y^{*}\}$, s={$s^{*}=\{s_{i}\}^{n}_{i=1}\oplus\{t^{*}_{i}\}^{m}_{i=1}\in\mathcal{S}^{*}$. $\oplus$ denotes the trigger injection operation, which could be identifier renaming (Sun et al., 2023; Li et al., 2024b; Yang et al., 2024) or dead-code insertion (Ramakrishnan and Albarghouthi, 2022; Wan et al., 2022; Li et al., 2024b). Subsequently, the attacker constructs the poisoned dataset $\mathcal{X}_{p}=\{\mathcal{X}\cup\mathcal{X}^{*}\}$ using the triggered samples. Finally, the model will be poisoned by training with $\mathcal{X}_{p}$ and minimizing the following loss function.
(2) $\footnotesize\begin{split}\mathcal{L}_{\mathcal{X}_{p}}\left(\theta^{*}\right)&=\underset{(s,y)\sim\mathcal{X}}{\mathbb{E}}\mathcal{L}\left(f\left(s;\theta^{*}\right),y\right)\\ &+\underset{\left(s^{*},y^{*}\right)\sim\mathcal{X}^{*}}{\mathbb{E}}\mathcal{L}\left(f\left(s^{*};\theta^{*}\right),y^{*}\right)\end{split}$
Where $\mathcal{L}(,)$ denotes the cross entropy loss. Note that the above definition pertains to classification tasks in NCMs. For another common code understanding task, the search task(e.g., code search), $s$ can be a text sequence, such as a query, and $y$ can be the ground-truth code. Therefore, the attacker first selects or inserts a query containing the target word as $\mathcal{X}^{*}$, then implants the trigger into the corresponding code as $y^{*}$, thereby constructing the poisoned sample $\mathcal{X}_{p}$. Then, the backdoor attack for search tasks also apply Equation 2 to train the model.
There are two types of trigger on backdoor attacks for code models. The first type is statement trigger backdoor where trigger is fixed or grammar dead code statement or snippet injected in code. The second type is identifier trigger where fixed or mixed tokens or words renaming the identifiers (function name/variables) in the code. Sun et al. (Sun et al., 2023) indicate that the token trigger is more stealthy than statement trigger. The statement trigger can be detect by human or static analysis tool easily. Therefore, we focus on the token trigger backdoor attacks which have the more serious threat.

### v2

# Eliminating Backdoors in Neural Code Models for Secure Code Understanding

@@ Abstract.
Neural code models (NCMs) have been widely used to address various code understanding tasks, such as defect detection. However, numerous recent studies reveal that such models are vulnerable to backdoor attacks. Backdoored NCMs function normally on normal/clean code snippets, but exhibit adversary-expected behavior on poisoned code snippets injected with the adversary-crafted trigger. It poses a significant security threat. For example, a backdoored defect detection model may misclassify user-submitted defective code as non-defective. If this insecure code is then integrated into critical systems, like autonomous driving systems, it could jeopardize life safety. Therefore, there is an urgent need for effective techniques to detect and eliminate backdoors stealthily implanted in NCMs.
To address this issue, in this paper, we innovatively propose a backdoor elimination technique for secure code understanding, called EliBadCode. EliBadCode eliminates backdoors in NCMs by inverting/reverse-engineering and unlearning backdoor triggers. Specifically, EliBadCode first filters the model vocabulary for trigger tokens based on the naming conventions of specific programming languages to reduce the trigger search space and cost. Then, EliBadCode introduces a sample-specific trigger position identification method, which can reduce the interference of non-backdoor (adversarial) perturbations for subsequent trigger inversion, thereby producing effective inverted backdoor triggers efficiently. Backdoor triggers can be viewed as backdoor (adversarial) perturbations. Subsequently, EliBadCode employs a Greedy Coordinate Gradient algorithm to optimize the inverted trigger and designs a trigger anchoring method to purify the inverted trigger. Finally, EliBadCode eliminates backdoors through model unlearning. We evaluate the effectiveness of EliBadCode in eliminating backdoors implanted in multiple NCMs used for three safety-critical code understanding tasks. The results demonstrate that EliBadCode can effectively eliminate backdoors while having minimal adverse effects on the normal functionality of the model. For instance, on defect detection tasks, EliBadCode substantially decreases the average Attack Success Rate (ASR) of the advanced backdoor attack from 99.76% to 2.64%, significantly outperforming the three baselines. The clean model produced by EliBadCode exhibits an average decrease in defect prediction accuracy of only 0.01% (the same as the baseline).
@@
copyright: rightsretaineddoi: 10.1145/3715782journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE063publicationmonth: 7ccs: Software and its engineering Software maintenance toolsccs: Security and privacy Software and application security
@@ 1. Introduction
Over the past decade, deep learning (DL)-based neural code models (NCMs) have demonstrated continuous improvement and impressive performance in handling software engineering (SE) tasks, particularly in code understanding tasks, such as defect detection (Wang et al., 2016; Zhou et al., 2019), code clone detection (Wei and Li, 2017; Fang et al., 2020), and code search (Sun et al., 2022, 2024a). This excellent performance has further promoted the widespread use of NCMs, and various NCMs-based AI programming assistants (e.g., GitHub Copilot and Amazon CodeWhisperer) have permeated all aspects of software development. Therefore, ensuring the security of NCMs is of paramount importance.
@@
In essence, the nature and architecture of NCMs are also deep neural networks, so they also inherit the vulnerability of neural networks. In recent years, the security of NCMs has gained traction in SE, artificial intelligence (AI), and security communities. Several existing works (Chen et al., 2024; Wan et al., 2022; Sun et al., 2023; Yang et al., 2024b; Li et al., 2024b; Su et al., 2025) have revealed that NCMs are vulnerable to a security threat called backdoor attacks. Such attacks, also called trojan attacks (Liu et al., 2018), aim to inject a backdoor pattern into the learned model with the malicious intent of manipulating the model’s outputs (Chen et al., 2024; Li et al., 2024a). Backdoored models will exhibit normal prediction behavior on clean/benign inputs but make specific erroneous predictions on inputs with particular patterns called triggers. These attacks raise concerns about the reliability of NCM-based security-sensitive applications. For example, the work (Sun et al., 2023) proposes a stealthy backdoor attack BadCode against NCMs for code search tasks. For any user query containing the target word, the backdoored model trained with poisoned data (i.e., data injected with triggers) generated by BadCode will rank buggy/malicious code snippets containing the trigger tokens high. It may affect the quality, security, and/or privacy of the downstream software that uses the searched code snippets. Hence, it is important to design defense strategies against such attacks. Currently, most backdoor defenses for NCMs are input detection defenses (Ramakrishnan and Albarghouthi, 2022; Hussain et al., 2023), which focus on detecting trigger-injected inputs to prevent the activation of backdoors in NCMs. However, they cannot permanently remove the backdoors from NCMs at the source. Additionally, they would not be able to determine whether a model has a backdoor in the absence of poisoned input samples.
@@
To address these issues, in this paper, we propose a novel backdoor defense technique named EliBadCode to eliminate backdoors in NCMs for secure code understanding. Specifically, EliBadCode first inverts (also called reverse engineers (Wang et al., 2019)) the attacker-crafted backdoor triggers from the backdoored NCM using a small number of available clean samples. This process is known as trigger inversion. Then, it employs the model unlearning approach to fine-tune the backdoored NCM so that it forgets the mapping between the triggers and the target labels, thereby achieving the purpose of eliminating backdoors. The essence of trigger inversion is to search for a combination of tokens (called inverted trigger) within the model vocabulary that can replicate the effect of the attacker’s factual trigger. To automate the search, EliBadCode transforms the trigger search into an optimization problem, where the inverted trigger is randomly initialized and iteratively updated using the Greedy Coordinate Gradient (GCG) algorithm (Zou et al., 2023). Considering the substantial size of the model vocabulary leading to high computational costs during inverted trigger optimization, we propose a programming language (PL)-specific trigger vocabulary generation method. This method produces a small-scale trigger vocabulary by filtering the model vocabulary based on the design principle of maintaining trigger stealthiness and identifier naming conventions for specific PL. Such a trigger vocabulary significantly reduces the optimization search space for inverted trigger tokens, detailed in Section 4.2. In addition, given that the GCG algorithm is prone to inverting non-backdoor (adversarial) perturbations at sensitive positions of the code, we propose a sample-specific trigger injection position identification method. It enables EliBadCode to inject the trigger into insensitive identifier positions for inverting, reducing the probability of inverting non-backdoor perturbations rather than effective triggers (which can also be viewed as backdoor perturbations), detailed in Section 4.3. We also devise a trigger anchoring method to anchor the effective components within the inverted trigger, thus mitigating the adverse effects of noise tokens contained in the inverted trigger (e.g., compromising the model’s normal performance). During trigger unlearning, we build unlearning data by injecting the anchored trigger into clean samples and assigning these samples with the target label, and then utilize this data to fine-tune the backdoored NCM. By controlling the trigger injection rate and the range of model parameter updating, EliBadCode can remove backdoors without affecting the normal performance of the model.
To evaluate the effectiveness of EliBadCode, we conduct comprehensive experiments, which involve three advanced backdoor attacks: CodePoisoner (Li et al., 2024b), BadCode (Sun et al., 2023), and AFRAIDOOR (Yang et al., 2024b), three code understanding tasks: defect detection, clone detection, and code search, and three model architectures: CodeBERT, CodeT5, and UniXcoder, a total of 27 attack scenarios. The results demonstrate that EliBadCode can significantly reduce the attack success rate (ASR) while maintaining nearly the same level of model prediction accuracy. For example, on defect detection tasks, EliBadCode can reduce the average ASR of the advanced attack BadCode from 99.76% to 2.64% with only 0.01% accuracy degradation on average, and is significantly better than three baselines ONION (Qi et al., 2021), DBS (Shen et al., 2022), and AttDef (Li et al., 2023c). In addition, we validate EliBadCode’s ability to eliminate backdoors in code large language models (LLMs). Specifically, we transfer the attack CodePoisoner to a popular code LLM called StarCoder (Li et al., 2023a), and then apply EliBadCode to eliminate backdoors in it. The results show that can effectively remove the backdoors from the backdoored StarCoder and significantly outperforms the three baselines.
@@
    We introduce two effective designs to reduce the cost of trigger inversion: PL-specific trigger vocabulary generation and sample-specific trigger injection position identification. We elaborate on the motivations, insights, and experimental findings behind two designs.
@@
    We evaluate the effectiveness of EliBadCode against three backdoor attacks (27 attack scenarios in total). The results demonstrate that EliBadCode can significantly reduce the ASR while maintaining the normal prediction accuracy of NCMs. We also validate EliBadCode’s ability to eliminate backdoors in code LLMs.
@@
    To the best of our knowledge, apart from EliBadCode, there are currently no dedicated techniques available for eliminating backdoors in NCMs. To foster advancement in this field and facilitate future researchers to verify, compare, and extend EliBadCode, we make the implementation code of EliBadCode (Sun et al., 2025) publicly available.
@@ 2.1. Code Understanding
Code understanding is a challenging task. Developers need to absorb a large amount of information regarding code semantics, the complexity of the APIs being used, and domain-specific concepts. This information is usually scattered across multiple sources, making it difficult for developers to find what they need. With the success of DL techniques, NCMs have been widely used for successfully addressing various code understanding tasks such as defect detection (Wang et al., 2016; Zhou et al., 2019), clone detection (Wei and Li, 2017; Fang et al., 2020), and code search (Wan et al., 2019; Sun et al., 2022). Given an NCM $f_{\theta}$, parameterized by $\theta$ and a clean dataset $\mathcal{X}=\{\mathcal{S},\mathcal{Y}\}$, where $s=\{s_{i}\}_{i=1}^{n}\in\mathcal{S}$ is a code snippet containing $n$ tokens, $y\in\mathcal{Y}$ is the ground-truth label. The model for code understanding tasks aims to minimize the following training loss:
(1) $\mathcal{L}\big{(}\theta\big{)}=\underset{(s,y)\sim\mathcal{X}}{\mathbb{E}}{-y\log(f_{\theta}(s))},$
where $\mathcal{L}(\cdot)$ is the cross-entropy loss. Note that Equation (1) is a general definition for the training objective of code understanding, which is widely used in existing works (Wang et al., 2016; Gu et al., 2018; Fang et al., 2020).
In recent years, with the success of the pre-training fine-tuning paradigm, a series of pre-trained models have been proposed to improve the performance of code understanding. Meanwhile, numerous studies demonstrate that these models face significant security threats, particularly backdoor attacks (Sun et al., 2023; Li et al., 2023b, 2024b). In this paper, we select the most representative pre-trained NCMs as the defense targets to eliminate backdoors, including CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021), and UniXcoder (Guo et al., 2022).
@@ 2.2. Backdoor Attacks
A backdoor attack can be defined as an attacker using hidden patterns to train a model, which produces the attacker’s specified output only when a specific trigger is present in the input (Wang et al., 2019; Han et al., 2024). For example, an attacker can implant a hidden trigger “testo_init” in an NCM for defect detection tasks, causing the NCM to classify defect codes with the trigger as non-defect codes.
In the backdoor attack, the attacker aims to train an NCM $f_{\theta}$ associated with a trigger $t^{*}=\{t^{*}_{i}\}^{m}_{i=1}$ with $m$ tokens and a target label $y^{*}\in\mathcal{Y}$. Specifically, the attacker first implants the trigger to a small number of samples $\mathcal{X}^{*}$, where $\mathcal{X}^{*}=\{\mathcal{S}^{*},y^{*}\}$, s={$s^{*}=\{s_{i}\}^{n}_{i=1}\oplus\{t^{*}_{j}\}^{m}_{j=1}\in\mathcal{S}^{*}$. $\oplus$ denotes the trigger injection operation, which could be identifier renaming (Sun et al., 2023; Li et al., 2024b; Yang et al., 2024b) or dead-code insertion (Ramakrishnan and Albarghouthi, 2022; Wan et al., 2022; Li et al., 2024b). Subsequently, the attacker constructs the poisoned dataset $\mathcal{X}_{p}=\{\mathcal{X}\cup\mathcal{X}^{*}\}$ using the triggered samples. Finally, the model will be poisoned by training with $\mathcal{X}_{p}$ and minimizing the following loss function:
(2) $\begin{split}\mathcal{L}_{\mathcal{X}_{p}}\left(\theta^{*}\right)&=\underset{(s,y)\sim\mathcal{X}}{\mathbb{E}}\mathcal{L}\left(f_{\theta^{*}}\left(s\right),y\right)+\underset{\left(s^{*},y^{*}\right)\sim\mathcal{X}^{*}}{\mathbb{E}}\mathcal{L}\left(f_{\theta^{*}}\left(s^{*}\right),y^{*}\right),\end{split}$
where $\mathcal{L}(\cdot)$ denotes the cross entropy loss. Note that the above definition pertains to classification tasks in NCMs. For another common code understanding task, the search task (e.g., code search), $s$ can be a text sequence, such as a natural language query, and $y$ can be the ground-truth code. Therefore, the attacker first selects or inserts a query containing the target word as $\mathcal{X}^{*}$, then implants the trigger into the corresponding code snippet as $y^{*}$, thereby constructing the poisoned sample $\mathcal{X}_{p}$. Then, the backdoor attack for search tasks also applies Equation (2) to train the model.
There are two types of triggers commonly used by backdoor attacks against NCMs. The first type is a statement trigger backdoor where the trigger is the fixed or grammar dead code statement or snippet injected in code. The second type is the identifier trigger where fixed or mixed tokens or words rename the identifiers (function name/variables) in the code snippet. The study (Sun et al., 2023) indicates that the token trigger is more stealthy than the statement trigger. The statement trigger can be detected by human or static analysis tools easily. Therefore, we focus on backdoor attacks with the token trigger, which have a more serious threat.